find the greatest common factor of 20a^3b, 32a^2b^2, and 44ab^3
Answer by CharlesG2(828)

find the greatest common factor of 20a^3b, 32a^2b^2, and 44ab^3
20 * a^3 * b
32 * a^2 * b^2
44 * a * b^3
greatest common factor means the greatest expression that will go into all three of these
20 * a^3 * b = 2 * 2 * 5 * a * a * a * b
32 * a^2 * b^2 = 2 * 2 * 2 * 2 * 2 * a * a * b * b
44 * a * b^3 = 2 * 2 * 11 * a * b * b * b
2 * 2 * a * b = 4 * a * b = 4ab is greatest common factor